# Satisficing Regret Minimization in Bandits

### v2

@@ 6.2 Concave Bandits
Setup: In this case, we consider an instance with arm set $[0,1]$ and concave reward function $r(x)=1-16(x-0.25)^{2}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.3$; for the non-realizable case, we set the satisficing level $S=1.5$. For both cases, we compare our algorithm SELECT with the convex bandit algorithm introduced in Agarwal et al. (2011). For SELECT, we use the algorithm in Agarwal et al. (2011) as the learning oracle. We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm space uniformly with stepsize $L^{-1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm space. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case are provided in Figure 3(a), and the results of non-realizable case are provided in Figure 3(b). From Figure 3(a), we can see that SELECT converges to satisficing arms faster and incurs smaller satisficing regret than the algorithm in Agarwal et al. (2011), SAT-UCB and SAT-UCB+. The reason why SAT-UCB and SAT-UCB+ has not converged is that they usually repeatedly pull non-satisficing arms which are close to satisficing level but provide empirically highest UCB. As shown in Figure 3(b), SELECT can also reach good enough distribution compared to the algorithm in Agarwal et al. (2011) , SAT-UCB, and SAT-UCB+ in the non-realizable case.
@@ 6.3 Lipschitz Bandits
Setup: In this case, we consider a two-dimensional Lipschitz bandit in domain $(x,y)\in[0,1]^{2}$ with Lipschitz rewards $f(x,y)=\min\{1,3e^{-100((x-0.5)^{2}+(y-0.7)^{2})}\}$. The length of the time horizon is set to be $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level S = 0.5; for the non-realizable case, we set the satisficing level S = 1.5. For both cases, we compare our algorithm SELECT with the uniformly discretized UCB introduced in Bubeck et al. (2011) (hereafter referred to as the “Uniform UCB”). For SELECT, we use the Uniform UCB as the learning oracle. We again use SAT-UCB and SAT-UCB+ as heuristics by uniformly discretizing the arm space with stepsize $L^{-1/4}T^{-1/4}$, where $L$ is the Lipschitz constant of the reward function, and run the algorithms over the discretized arms. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case are provided in Figure 4(a), and the results of non-realizable case are provided in Figure 4(b). From Figure 4(a), we can see that SELECT converges to satisficing arms faster and incurs smaller satisficing regret than Uniform UCB, SAT-UCB and SAT-UCB+. From 4(b), we can see that SELECT still works the best while Uniform UCB, SAT-UCB and SAT-UCB+ have similar performance.
@@ 7 Conclusion
In this paper, we propose SELECT, a general algorithmic template, for satisficing regret minimization in bandit optimization. For a given bandit optimization problem and a sub-linear regret learning oracle, we show that SELECT attains constant satisficing regret in the realizable case and the same regret as the learning oracle in the non-realizable case. We instantiate SELECT to finite-armed bandits, concave bandits, and Lipschitz bandits, and also make a discussion on the corresponding lower bounds. Finally, numerical experiments show that SELECT attains constant regret in several popular bandit problems and demonstrates excellent performances when compared to other benchmarks in both the realizable and the non-realizable cases.
For future studies, we suggest further exploration in the following directions. First, our work mainly focuses on the stationary settings, and further research could explore satisficing in the nonstationary online learning settings (Wei and Luo 2021, Cheung et al. 2022). Also, the notion of satisficing regret indeed discourages exploration. Therefore, another direction is to study the role of satisficing regret in settings with limited adaptivity (Gao et al. 2019, Feng et al. 2024).
@@ Appendix A Proof of Theorem 1
where $M_{\beta}=\sum_{k=1}^{\infty}k^{\beta}\cdot 2^{-k}$, and the total satisficing regret is bounded by
$\texttt{Regret}_{S}\leq\dfrac{8L_{\alpha,\beta}^{\alpha/(1-\alpha)}C_{1}^{1/(1-\alpha)}(M_{\beta}+2)}{(1-\alpha)^{\beta}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\alpha/(1-\alpha)}\log\left(\dfrac{D_{\alpha,\beta}C_{1}}{\Delta_{S}^{*}}\right)^{\alpha\beta^{\prime}/(1-\alpha)+1}.$
@@
$\texttt{Regret}_{S}\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}$
@@ A.1 Proof of Proposition 1
$\mathbb{E}\left[\sum_{s=1}^{t_{i}}\max\{S-r(X_{s}),0\}\right]\leq\mathbb{E}\left[t_{i}r(X^{*})-\sum_{s=1}^{t_{i}}r(X_{s})\right]\leq$ $C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}$
@@
Since $\hat{X}_{i}$ is selected randomly according to the trajectory of $\texttt{ALG}(t_{i})$ in round $i$, we further have
$\mathbb{E}[\max\{S-r(\hat{X}_{i}),0\}]\leq\mathbb{E}[r(X^{*})-r(\hat{X}_{i})]=$ $\frac{\mathbb{E}\left[t_{i}r(X^{*})-\sum_{s=1}^{t_{i}}r(X_{s})\right]}{t_{i}}$
@@
Therefore the satisficing regret incurred when pulling $\hat{X}_{i}$ for $T_{i}=\lceil\gamma_{i}^{-2}\rceil$ times is bounded by
@@
If $r(\hat{X}_{i})<S$, then for every $k$, round $i$ is not terminated after pulling arm $\hat{X}_{i}$ for $k$ times (in addition to the first $T_{i}$ times) occurs with probability at most
@@
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>r(\hat{X}_{i})\right)$
@@
Therefore expected number of arm pulls of $\hat{X}_{i}$ after first pulling it $T_{i}$ times is bounded by
$\mathbb{E}\left[\text{Number of times }\hat{X}_{i}\text{ is pulled in round }i\text{ after the first }T_{i}\text{ times}\right]$
@@ A.2 Proof of Proposition 2
$\mathbb{E}[r(X^{*})-r(\hat{X}_{i})]=$ $\dfrac{1}{t_{i}}\mathbb{E}\left[t_{i}r(X^{*})-\sum_{s=1}^{t_{i}}r(X_{s})\right]\leq\dfrac{1}{t_{i}}C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}=C_{1}t_{i}^{\alpha-1}\log(t_{i})^{\beta}$
@@
$\text{Pr}(r(X^{*})-r(\hat{X}_{i})\geq\Delta_{S}^{*}/2)\leq\text{Pr}\left(r(X^{*})-r(\hat{X}_{i})\geq\dfrac{16C_{1}}{(1-\alpha)^{\beta}}\log(2/\gamma_{i})^{\beta}\right)\leq\dfrac{1}{8}.$ (4)
Now, if $r(\hat{X}_{i})\geq r(X^{*})-\Delta_{S}^{*}/2$, which happens with probability at least 7/8 according to the above inequality, then $r(\hat{X}_{i})\geq S+\Delta_{S}^{*}/2$ holds. On the other hand, if $i$ satisfies the condition stated in the proposition, then $\gamma_{i}\log(1/\gamma_{i})^{1/2}\leq\Delta_{S}^{*}/16$.
Note that $f(x)=\log(x)/x$ is monotonically decreasing when $x\geq 3$ and $32\cdot\log(2)/2\geq\log(3)/$ 3, for all $k\geq 0$,
@@
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}<S=r(X^{*})-\Delta_{S}^{*}\leq r(\hat{X}_{i})-\dfrac{\Delta_{S}^{*}}{2}\leq r(\hat{X}_{i})-2\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}},$
@@
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}\leq r(\hat{X}_{i})-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}.$
@@
$\text{Pr}\left(\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}\leq r(\hat{X}_{i})-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}\right)\leq\dfrac{1}{(T_{i}+k)^{2}}\leq\dfrac{1}{(k+4)^{2}}.$
Therefore conditioned on $r(\hat{X}_{i})\geq r(X^{*})-\Delta_{S}^{*}/2$ the probability that round $i$ is terminated by the end of the horizon is bounded by

### v3

@@ 7.2 Concave Bandits
Setup: In this case, we consider an instance with arm set $[0,1]$ and concave reward function $r(x)=1-16(x-0.25)^{2}$. We vary the length of the time horizon from $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.3$; for the non-realizable case, we set the satisficing level $S=1.5$. For both cases, we compare our algorithm SELECT with the convex bandit algorithm introduced in Agarwal et al. (2011). For SELECT, we use the algorithm in Agarwal et al. (2011) as the learning oracle. We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm set uniformly with stepsize $L^{-1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm set. The experiment is repeated for 1000 times, and we report the average results.
Results: The results of realizable case are provided in Figure 3(a), and the results of non-realizable case are provided in Figure 3(b). One can see from Figure 3(a) that the expected satisficing regret of SELECT becomes stable after $T$ exceeds $500$, suggesting that SELECT attains a constant expected satisficing regret in the realizable case. On the other hand, all the other three algorithms fail to exhibit a constant expected satisficing regret in the realizable case. When comparing the expected satisficing regret of the four algorithms, we observe that SELECT incurs a smaller expected satisficing regret than all three other algorithms.
One can see from Figure 3(b) that in the non-realizable case, SELECT incurs less expected standard regret compared to the algorithm in Agarwal et al. (2011). Compared to SAT-UCB and SAT-UCB+, the expected standard regret of SELECT is larger, but the performance of SELECT is comparable with both SAT-UCB and SAT-UCB+.
In this experiment, the reason why SELECT outperforms the algorithm in Agarwal et al. (2011) can be explained as follows. In the instance we use in this experiment, both the set of optimal arms and the set of significantly suboptimal arms are large, and directly running the algorithm in Agarwal et al. (2011) over the entire time horizon tend to over-explore due to larger confidence intervals, thus would potentially incur significant regret. In contrast, SELECT starts with running the algorithm in Agarwal et al. (2011) over a small number of time steps $t_{i}$. This allows SELECT to have a much smaller confidence intervals in early stages of the algorithm, thus explores less and exploits more aggressively. Furthermore, when running forced sampling in Step 2, SELECT also exploits a reasonably good arm obtained from running the learning oracle, which means even more exploitation for SELECT. Therefore, although theoretically SELECT may underperform Agarwal et al. (2011) by a logarithmic factor of $T$ in the worst case, in the instance used in Figure 3(b), SELECT outperforms the algorithm in Agarwal et al. (2011). Similar phenomenon can also be observed in the numerical experiment for Lipschitz bandits in Section 7.3.
@@ 7.3 Lipschitz Bandits
Setup: In this case, we consider a two-dimensional Lipschitz bandit in domain $(x,y)\in[0,1]^{2}$ with Lipschitz rewards $f(x,y)=\min\{1,3e^{-100((x-0.5)^{2}+(y-0.7)^{2})}\}$. We vary the length of the time horizon from $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level S = 0.5; for the non-realizable case, we set the satisficing level S = 1.5. For both cases, we compare our algorithm SELECT with the uniformly discretized UCB introduced in Bubeck et al. (2011) (hereafter referred to as the “Uniform UCB”). For SELECT, we use the Uniform UCB as the learning oracle. We again use SAT-UCB and SAT-UCB+ as heuristics by uniformly discretizing the arm set with stepsize $L^{-1/4}T^{-1/4}$, where $L$ is the Lipschitz constant of the reward function, and run the algorithms over the discretized arms. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case are provided in Figure 4(a), and the results of non-realizable case are provided in Figure 4(b). One can see from Figure 4(a) that the expected satisficing regret of SELECT becomes stable after $T$ exceeds $500$, suggesting that SELECT attains a constant expected satisficing regret in the realizable case. On the other hand, none of the other three algorithms are able to attain a constant expected satisficing regret. When comparing the expected satisficing regret of the four algorithms, one can observe that SELECT incurs an expected satisficing regret smaller than all three other algorithms.
As for the non-realizable case, one can see from Figure 4(b) that SELECTachieves the best empirical performance among the four algorithms, while the empirical performance of SAT-UCB and SAT-UCB+ is similar to that of Uniform UCB in the non-realizable case.
@@
7.4 Tail Risk Control in Finite-Armed Bandits
We test the performance of SELECT-LITE in terms of both expected satisficing regret and tail distribution of satisficing regret using an instance of finite-armed bandit.
Setup: We consider an instance of 4 arms. The expected rewards of all arms are set as $\{0.2,0.4,0.6,0.8\}$. For the realizable case, we set the satisficing level $S=0.7$; for the non-realizable case, we set the satisficing level $S=1.5$. For SELECT and SELECT-LITE, we use Thompson sampling as the learning oracle. We vary $\zeta\in\{0.1,0.4\}$.
We first compare the expected satisficing regret of SELECT-LITE with that of SELECT and Thompson sampling. In this case, we vary the length of the time horizon from 1000 to 10000 with a stepsize of 1000. The experiment is repeated for 1000 times and we report the average satisficing regret. Then we compare the empirical tail distribution of satisficing regret between SELECT-LITE and SELECT under time horizon $T=10000$. The experiment is repeated for 10000 times.
Results: The results of realized satisficing regret distribution under $\zeta=0.1$ are presented in Figure 5(a) for SELECT and Figure 5(b) for SELECT-LITE , and the comparison on tail probability is presented in Figure 5(c). The results under $\zeta=0.4$ are presented in Figure 6. Here we report the empirical distribution of realized satisficing regret using histograms with a bin width of 10. From Figure 5(c) and 6(c), one can see that under both $\zeta=0.1$ and $\zeta=0.4$, SELECT-LITE exhibits a lower tail probability compared to SELECT. When comparing between between the case of $\zeta=0.1$ and the case of $\zeta=0.4$, SELECT-LITE with $\zeta=0.1$ exhibits a noticeably thinner tail, even though a larger $\zeta$ should theoretically yield the faster decay rate. This is because when $\zeta=0.1$, $t_{i}^{\prime}$ increases rapidly in the first few rounds, thus SELECT-LITE can quickly find the satisficing arm within a few rounds. Consequently, the empirical tail of SELECT-LITE when $\zeta=0.4$ is a summation of tails of many rounds, but the one when $\zeta=0.1$ is a summation of simply one or two rounds, which potentially makes the tail probability smaller.
The results of expected satisficing regret in the realizable case are presented in Figure 7(a). The results of expected standard regret in the non-realiable case are presented in Figure 7(b). From Figure 7(a) one can observe that SELECT-LITE  exhibits constant satisficing regret that is only slightly larger than that of SELECT. Furthermore, both when $\zeta=0.1$ and when $\zeta=0.4$, SELECT-LITE incur less satisficing regret compared to Thompson sampling. From Figure 7(b) one can see that SELECT-LITE  with different $\zeta$ incurs roughly the same standard regret as SELECT and is slightly better than Thompson sampling. This is because each SELECT-LITE algorithm shares a backbone with SELECT, and SELECT-LITE with $\zeta=0.1$ has a $O(T^{0.55})$ regret bound that is very close to that of SELECT. For SELECT-LITE with $\zeta=0.4$, although the theoretical standard regret is $O(T^{0.7})$, worse than the $O(T^{0.5})$ guarantee for Thompson sampling, it outperforms Thompson sampling empirically because its aggressive exploration yields a sizable constant-factor advantage over the time horizons studied. With these observaations, we conclude that by running SELECT-LITE instead of SELECT, the decision maker can have better protection against tail risks, both theoretically and empirically, without incurring a significantly larger satisficing regret in expectation.
(a) SELECT Regret Distribution
(b) SELECT-LITE Regret Distribution
(c) Tail Comparison
Figure 5: Realized satisficing regret distribution of SELECT  and SELECT-LITE  at $T=10000$ with $\zeta=0.1$
(a) SELECT Regret Distribution
(b) SELECT-LITE Regret Distribution
(c) Tail Comparison
Figure 6: Realized satisficing regret distribution of SELECT  and SELECT-LITE  at $T=10000$ with $\zeta=0.4$
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 7: Comparison of SELECT , SELECT-LITE  with different $\zeta$, and Thompson sampling
8 Numerical Experiments on Real Data: Satisficing in Dynamic Pricing
To further evaluate the empirical performance of our algorithm in more practical settings, in this section, we conduct numerical experiments on an instance calibrated from the avocado dataset Kiggins (2018).
8.1 Data Description, Pre-Processing, and Setup
We use the avocado dataset Kiggins (2018) in the New York region. The data records weekly average prices of organic avocado and the corresponding weekly total selling volume from January 4, 2015, to December 27, 2015, for a total of 52 weeks (with outliers deleted). The weekly average prices in these 52 data points range from \$1.65 to \$2.34, and the average price is \$2.03. The weekly selling volumes range from 9205 to 25675, and the average weekly selling volume is 17106.
We consider a dynamic pricing problem with linear demands to the prices, and test the performance of the algorithms on this model fitted by the avocado dataset. Assuming that the demand is equal to $D_{t}=g-hp_{t}+\epsilon_{t}$, where $\epsilon_{t}\sim N(0,\sigma^{2})$, given price $p_{t}$ at each time step $t$, we fit the linear regression model from the real data to be $\hat{g}=32724,\;\hat{h}=7678,\;\hat{\sigma}=4100.$ By normalizing the model to ensure unit noise variance, we obtain the coefficients to be
$\hat{g}_{n}=32724/4100,\;\hat{h}_{n}=7678/4100,\;\hat{\sigma}_{n}=1.$
We also extend the range of average price to be $p\in[0,4]$. The mean revenue received at each time $t$ is $p_{t}(\hat{g}_{n}-\hat{h}_{n}p_{t})$, and the mean revenue of each feasible price is visualized in Figure 8.
Figure 8: Expected revenue of the feasible prices $p\in[0,4]$
We set the length of the time horizon to be 500 to 8000 for the realizable case and 5000 for the non-realizable case with a stepsize of 500. For the realizable case, we set the satisficing level $S=8.0$; for the non-realizable case, we set the satisficing level $S=10.0$. For both cases, we compare our algorithm SELECT with Linear UCB introduced in Lattimore and Szepesvári (2020). For SELECT, we use Linear UCB in Lattimore and Szepesvári (2020) as the learning oracle. We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm set uniformly with a stepsize $L^{-1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm set. The experiment is repeated 1000 times and we report the averaged results.
8.2 Results
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 9: Comparison of  SELECT , Linear UCB, SAT-UCB and SAT-UCB+ for linear bandit
The results of the realizable case are provided in Figure 9(a), and the results of the non-realizable case are provided in Figure 9(b). From Figure 9(a), one can observe that the expected satisficing regret of SELECT becomes stable after $T$ exceeds $3000$. This suggests that in the realizable case, SELECT is able to attain constant expected satisficing regret. On the other hand, none of the other three algorithms are able to achieve constant expected satisficing regret in the realizable case. When comparing the expected satisficing regret of the four algorithms, one can see that SELECT incurs a smaller expected satisficing regret than all the other three algorithms.
From Figure 9(b), one can see that in the non-realizable case, SELECT slightly outperforms Linear UCB and significantly outperforms both SAT-UCB and SAT-UCB+. The reason why SELECT and Linear UCB significantly outperforms SAT-UCB and SAT-UCB+ is that in the problem, SAT-UCB and SAT-UCB+ views the problem as a Lipschitz bandit, without making use of the linear structure of demand.
@@ 9 Conclusion
In this paper, we propose SELECT, a general algorithmic template, for satisficing regret minimization in bandit optimization. For a given class of bandit optimization problems and a sub-linear regret learning oracle, we show that SELECT attains constant expected satisficing regret in the realizable case and the same regret as the learning oracle in the non-realizable case. We instantiate SELECT to finite-armed bandits, concave bandits, and Lipschitz bandits, and also make a discussion on the corresponding lower bounds. We also consider tail risk control in satisficing regret minimization, where we introduce SELECT-LITE, which attains a light-tailed satisficing regret distribution while still maintaining a constant expected satisficing regret. Finally, we conduct numerical experiments to validate the performance of SELECT and SELECT-LITE on both synthetic datasets and a real-world dynamic pricing case study.
For future studies, we suggest further exploration in the following directions. First, our work mainly focuses on the stationary settings, and further research could explore satisficing in the nonstationary online learning settings (Wei and Luo 2021, Cheung et al. 2022). Also, the notion of satisficing regret indeed discourages exploration. Therefore, another direction is to study the role of satisficing regret in settings with limited adaptivity (Gao et al. 2019, Feng et al. 2023).
@@ Appendix A Proof of Theorem 1
where $M_{\beta}=\sum_{k=1}^{\infty}k^{\beta}\cdot 2^{-k}$ is a constant only dependent on $\beta$, and the total satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\dfrac{8L_{\alpha,\beta}^{\alpha/(1-\alpha)}C_{1}^{1/(1-\alpha)}(M_{\beta}+2)}{(1-\alpha)^{\beta}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\alpha/(1-\alpha)}\log\left(\dfrac{D_{\alpha,\beta}C_{1}}{\Delta_{S}^{*}}\right)^{\alpha\beta^{\prime}/(1-\alpha)+1}.$
@@
$\mathbb{E}[\texttt{Regret}_{S}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}$
@@ A.1 Proof of Proposition 1
$\mathbb{E}\left[\sum_{s=1}^{t_{i}}\max\{S-r(A_{s}),0\}\right]\leq\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq$ $C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}$
@@
Since $\hat{A}_{i}$ is selected randomly according to the trajectory of $\texttt{ALG}(t_{i})$ in round $i$, we further have
$\mathbb{E}[\max\{S-r(\hat{A}_{i}),0\}]\leq\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]=$ $\frac{\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]}{t_{i}}$
@@
Therefore the satisficing regret incurred when pulling $\hat{A}_{i}$ for $T_{i}=\lceil\gamma_{i}^{-2}\rceil$ times is bounded by
@@
If $r(\hat{A}_{i})<S$, then for every $k$, round $i$ is not terminated after pulling arm $\hat{A}_{i}$ for $k$ times (in addition to the first $T_{i}$ times) occurs with probability at most
@@
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>r(\hat{A}_{i})\right)$
@@
Therefore expected number of arm pulls of $\hat{A}_{i}$ after first pulling it $T_{i}$ times is bounded by
$\mathbb{E}\left[\text{Number of times }\hat{A}_{i}\text{ is pulled in round }i\text{ after the first }T_{i}\text{ times}\right]$
@@ A.2 Proof of Proposition 2
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]=$ $\dfrac{1}{t_{i}}\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq\dfrac{1}{t_{i}}C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}=C_{1}t_{i}^{\alpha-1}\log(t_{i})^{\beta}$
@@
$\text{Pr}(r(A^{*})-r(\hat{A}_{i})\geq\Delta_{S}^{*}/2)\leq\text{Pr}\left(r(A^{*})-r(\hat{A}_{i})\geq\dfrac{16C_{1}}{(1-\alpha)^{\beta}}\log(2/\gamma_{i})^{\beta}\right)\leq\dfrac{1}{8}.$ (7)
Now, if $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$, which happens with probability at least 7/8 according to the above inequality, then $r(\hat{A}_{i})\geq S+\Delta_{S}^{*}/2$ holds. On the other hand, if $i$ satisfies the condition stated in the proposition, then $\gamma_{i}\log(1/\gamma_{i})^{1/2}\leq\Delta_{S}^{*}/16$.
Note that $f(y)=\log(y)/y$ is monotonically decreasing when $y\geq 3$ and $32\cdot\log(2)/2\geq\log(3)/$ 3, for all $k\geq 0$,
@@
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}<S=r(A^{*})-\Delta_{S}^{*}\leq r(\hat{A}_{i})-\dfrac{\Delta_{S}^{*}}{2}\leq r(\hat{A}_{i})-2\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}},$
@@
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}\leq r(\hat{A}_{i})-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}.$
@@
$\text{Pr}\left(\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}\leq r(\hat{A}_{i})-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}\right)\leq\dfrac{1}{(T_{i}+k)^{2}}\leq\dfrac{1}{(k+4)^{2}}.$
Therefore conditioned on $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$ the probability that round $i$ is terminated by the end of the horizon is bounded by